OpenStudy (anonymous):

I have manually worked out the formula for compounding interest on $5,000 over 3 years @ 6% paid 6 monthly, but can anyone tell me how do do it on excel spreadsheet please.

OpenStudy (anonymous):

Hi All I think I have worked it out can someone confirm that it is the correct way to enter. =5000*(1+(0.06/2))^(3*2)

OpenStudy (amistre64):

well, that is the correct formula i would do: a b c d 1 0 5000 =b1*.06/2 =b1+c1 2 1 =d1 =b2*.06/2 =b2+c2 3 2 4 3 5 4 6 5